OpenStudy (callisto):

Uncategorised question #4 Let \[x= 0.\dot{1}\dot{7} + 0.0\dot{1}\dot{7} + 0.00\dot{1}\dot{7} +...\], find the value of x. * there are dots on the number 1 and 7 * Once again, no caluclator

OpenStudy (experimentx):

GS

OpenStudy (experimentx):

a = 0.17 r = 1/10

OpenStudy (callisto):

First term = 17/99 r = 1/10 Sum = 17/99 / (1-1/10) = 17/99 / (9/10) = 170/891 ?

OpenStudy (experimentx):

i thought first tern is 17/100

OpenStudy (experimentx):

nice technique ...

OpenStudy (callisto):

17/99 It's recurring decimal, but I don't know how to type it out :|

OpenStudy (callisto):

I got the answer though. Thanks! :) (The solution given by my teacher few years ago was wrong.. no wonder why I can't get his answer :| )

OpenStudy (experimentx):

lol .. you remember solution of few years ago ... you've got quite a memory!!

OpenStudy (callisto):

Not remember, I just marked down the final answer on the paper :|